A is the set of +ve integers such that when divide by 2, 3, 4, 5 adn 6 leaves the remainders 1, 2, 3, 4 and 5 respectively. How many integer(s) between 0 and 100 belongs to set A?

Grade:10

1 Answers

Srishti Nanduri
9 years ago
let a = 2q+1 a= 3q+2 a= 4q+3 a= 5q+4 a= 6q+5 Now in the equation a= 6q+5, Let q=1 than a = 6q +5 = 6*1 +5 = 11 If we replace a with eleven with ALL the above equations .. we can infer that 11 satisfies all the equations... So ELEVEN IS THE ONLY NO BETWEEN 0 AND 100 WHICH BELONGS TO SET A...... I hope it helps